OpenStudy (anonymous):

What is (-1)^(3/5) and (-1)^(6/10)?

OpenStudy (anonymous):

First, since 6/10 is reducible to 3/5, both the given expressions evaluate to the same value. To evaluate the expression, break down the exponent into two parts ^3 and ^(1/5). The first part is -1^3, which equals -1 * -1 * -1 = -1. The second part is -1^(1/5) = the fifth root of -1. Can you think of a number, which when multiplied with itself five times, gives you -1? That is the answer

OpenStudy (anonymous):

Can i write (-1)^(6/10) as( (-1)^6)^(1/10) which is 1?

OpenStudy (solomonzelman):

\[ (-1)^{3/5}~~~and~~(-1)^{6/10}\]same as \[ \sqrt[5]{(-1)^3}~~~and~~ \sqrt[10]{(-1)^6}\]which gives\[ \sqrt[5]{-1}~~~and~~ \sqrt[10]{1}\]\[ -1~~~and~~ 1\]

OpenStudy (solomonzelman):

Whatever "and" is supposed to mean here.

OpenStudy (anonymous):

@SolomonZelman I think it is (-1)^(3/5)

OpenStudy (kainui):

Technically the first one has 5 answers and the second one has 10 answers since there are n-roots. For instance, what's the square root of 4? There are two answers, -2 and +2. Similarly what are the fourth roots of 16? +2, -2, i2, and -i2. See? (i2)^4=16 The general formula for the n-roots of 1 is:\[e^{i 2\pi/n}=\cos(\frac{2\pi}{n})+i \sin(\frac{2\pi}{n})\]